### v1

# P.O. Box 15500, 00076 Aalto, Finland

Abstract
A key issue with standard network measures of closeness and betweenness centrality is that they rely on the shortest paths between nodes within the network structure, whereas the degree centrality only reveals the immediate neighborhood of a node. Furthermore, many measures found in the literature do not accurately represent the physical or probabilistic characteristics of nodal centrality, network flow, and other salient properties. For example, recurrent spreading in a network is often overlooked by these metrics. Standard centrality measures have limitations, being optimal for one application but not for others. Here, we present new metrics based on our influence spreading model to characterize network structure for various network science applications. These probabilistic measures account for all feasible walks and cycles in the network. We compare our new metrics with the standard metrics in terms of the node rankings given by different centrality measures, by examining scatter plots, and by using the Pearson correlation coefficient. In the influence spreading model, we define the in-centrality measure to characterize how central a node is as a target of influence by other nodes and the out-centrality measure to characterize how central a node is as a source of influence on other nodes. Our results show that the influence spreading betweenness centrality reveals the importance of alternative routes while maintaining similarity to standard betweenness centrality.
keywords:
centrality metrics, influence spreading model, in-centrality, out-centrality, centrality metric comparison, betweenness centrality
1 Introduction
Network science is a rapidly expanding field of research that focuses on investigating the structure and dynamics of interdependent phenomena in complex systems of various scientific disciplines. This approach has been applied especially in studies of social science and communication networks, but also in studies of biological networks and its numerous subfields. In these fields of research, various modeling and other computational methodologies have become important in characterizing and analyzing the properties of these networks. The methods used in network science comprise a range of approaches, including statistical techniques, visualization methods, and various metrics that describe the structure, properties, and processes in networks. An important area of research focuses on clusters and communities in networks, where numerous methods for detecting communities have been proposed. In addition, basic tools, such as network centrality metrics, help measure the importance of individual nodes within the network structure.
In this article, we will frequently refer to paths, walks, and cycles. Based on [bender2010] and [bovet2022centralities], we define a walk on a network as an alternating sequence of nodes and edges, where every node is incident to both the edges that come before and after it in the sequence, and a path on a network is a walk where all nodes and edges are distinct. It is important to emphasize that all paths are walks, but all walks are not paths. A cycle is a walk where edges are required to be distinct, and the only repeated nodes are the first and the last node of the walk [bender2010].
Standard network centrality metrics, such as degree centrality, closeness centrality, and betweenness centrality, have become common in network science. The degree centrality $C_{D}(v)$ of node $v$ measures the number of neighbors $k_{v}$ that the node has with respect to the total possible number of neighbors [saxena2020centrality], as follows
$C_{D}(v)=\frac{k_{v}}{N-1},$ (1)
where $N$ is the total number of nodes in the network. The closeness centrality $C_{C}(v)$ of a node $v$ measures how close the node is to all other nodes in the network [saxena2020centrality]. It is calculated as the reciprocal of the average length of the shortest paths between node $v$ and all other nodes in the graph:
$C_{C}(v)=\frac{1}{\frac{\sum_{u\neq v}{d(u,v)}}{N-1}}=\frac{N-1}{\sum_{u\neq v}{d(u,v)}}.$ (2)
The betweenness centrality $b(v)$ measures how frequently a node $v$ lies on the shortest path between two other nodes in the network:
$b(v)=\sum_{\begin{subarray}{c}s\neq v\neq t\\ s\neq t\end{subarray}}\frac{\sigma_{st}(v)}{\sigma_{st}},$ (3)
where $\sigma_{st}$ represents the total number of shortest paths from node $s$ to node $t$, while $\sigma_{st}(v)$ denotes the number of paths that pass through node $v$  [freeman1977set]. The high value of betweenness centrality indicates that a node often acts as a gatekeeper of information or resources. Numerous variants of both the closeness and betweenness centrality can be found in the literature [saxena2020centrality, bockholt2018, bockholt2021].
It should be noted that standard centrality measures have limitations, as they may be optimal for one application but suboptimal for another. In order to choose metrics, characteristics to consider are whether the network is directed or undirected and what is being analyzed, such as nodes with high degree (hubs) or those with mediating control, influential nodes, or the smallest set of nodes that can spread information [bovet2022centralities]. Moreover, it is important to assess variations in the ranking of different centrality measures to understand their contextual importance.
In the present study and being inspired by the influence spreading model [kuikkaSciRep], we propose new metrics to describe consistent network characteristics, which are designed to be applicable in a wide range of network science scenarios. These new metrics consider all feasible paths, walks, and cycles within the network structure. In Section 2, we discuss how the influence spreading model is related to several important characteristics of networks. In Section 3, we give an overview of the application areas of the influence spreading model. In Section 4, we briefly discuss the research related to our study. In Section 5, we introduce the influence spreading model, give a motivating example of how the model works, and finally introduce our new metrics. In Section 6, we present the results of how the new metrics behave on multiple data sets compared to standard centrality metrics. In Section 7, we discuss our research results and how they relate to network modeling. Finally, in Section 8 we summarize our findings.
2 Network Structure and Dynamics
Here we highlight key characteristics of networks with new metrics that describe their structures, focusing on nodes and edges [kuikkaSciRep]. Additionally, we describe network flow as a spreading process of non-exclusive events within the network structure.
In the influence spreading model, weights are assigned to both nodes and edges. In this probabilistic framework, the weight of an edge represents the probability of influence or information passing through that edge, while the weight of a node indicates the probability of the node changing its state and enabling the passage of influence or information. These two properties can also be modeled as distinct features.
In the present study, the edges are considered to be directed, unidirectional, or bidirectional. For the bidirectional edges, it is possible to assign different weights for each direction. Undirected edges can be modeled by setting the same edge weight to each direction in a bidirectional edge. Then removing an edge or node from the network structure is equivalent to setting its weight to zero.
One of the most important features of our model is that cycles (or loops) [jiang2025identifying] are allowed in the network structure without restrictions. Moreover, these cycles can exist within other cycles and bidirectional links can be included in cycles, allowing flow to occur in both directions. An example of this is brain networks, where the cyclic propagation of signals can take place.
The definitions of standard network centrality measures, such as closeness centrality and betweenness centrality, count only geodesic or shortest paths between two nodes in the network. This is a limitation because it overlooks other paths and walks that can significantly influence various centrality properties and flow characteristics in the network. Our proposed centrality metrics address this issue by considering all possible walks that are shorter than a maximum walk length, which is defined as a model parameter. When the edge probabilities are low, due to decay effects, the model accounts for all relevant walks with the accuracy of the computations.
Centrality measures inherently imply specific patterns of flow in networks [borgatti2005]. This is true even for degree centrality, unless the network graph is used solely for visualization purposes. Two main extreme types of processes occur in networks: spreading along self-avoiding walks and spreading along node-state-independent walks. In the case of self-avoiding walks, each node can only be visited once, whereas in the case of node-state-independent walks, there are no such restrictions. Evidently, cycles cannot appear on self-avoiding walks.
Network centrality metrics and community detection methods have been developed both for local and global network structures  [qiu2021ranking]. Here the term global refers to approaches that take into account the entire network. In many cases, global methods address the limitations found in standard metrics or other measures proposed in the literature. Many of them are based on existing techniques, enhance them with improvements or modifications, or use combinations of standard metrics [qiu2021ranking, meng2025development, kouam2024novel].
3 Applications
In this section, we highlight some representative applications where our proposed metrics play a significant role. This list is not exhaustive, as network centrality metrics are crucial in most network analysis studies.
Social network analysis can help us find influential leaders in a social group, trace the source of new ideas and innovations, and assess the speed of information spreading [brass2022]. In social network analysis, an important application is the detection of the source or sources of rumor, misinformation, and disinformation [shelke2019]. The characteristics of influence spreading involve complex social contagion mechanisms, where multiple interactions are necessary for an opinion or behavior change to occur [centola2007]. In contrast, information or persuasive comments may require just a single interaction to propagate the news. It is important to note that we use the term complex social contagion to refer to Damon Centola’s concept of complex contagion [centola2007], where multiple influences are needed for an individual to change their opinion or behavior. In section 5.1, we describe in detail a different meaning for the term complex contagion in the context of the influence spreading model.
Modeling the spread of epidemics has become a significant area of research due to the COVID-19 pandemic. New approaches have been developed to describe both the spread and containment of the virus. Detailed modeling at the organizational level has allowed for targeted isolation of specific units within organizations and has enabled the investigation of how individuals in different contact networks contribute to spreading the virus [kuikkaSciRep]. Commonly used metrics to summarize model outcomes include network centrality, betweenness, and the basic reproduction number.
Communication networks are usually evaluated using various metrics that indicate their connectivity and reliability. In this context, circular propagation is not relevant, which allows for the use of self-avoiding walk models. When investigating cyber threats and attacks, it is important to model at both the enterprise and cross-enterprise levels. The use of various technical attack techniques and multi-layer approaches highlights the necessity for detailed modeling of these environments [kuikka2025network].
4 Related Work
Centrality measures based on shortest paths make simplistic assumptions of how information, traffic or other network processes flow through a network [freeman1991, borgatti2005, bockholt2021]. The spread of information can either randomly or by intention take a longer and indirect route with many intermediaries [freeman1991]. Stephen Borgatti noted how a flow process can use unrestricted walks and where there can either be transfer (no duplication), serial duplication or parallel duplication depending on what the flow process consists of [borgatti2005]. Bockholt and Zweig noted that the use of certain centrality measures contains an implicit assumption on how the network process flows through the network [bockholt2018]. In another publication, the authors also note that standard betweenness centrality is not the best at identifying influential nodes in an information spreading process [bockholt2021]. However, Freeman [freeman1978] and Krackhardt [krackhardt1990] argued that standard betweenness centrality is closely associated with social power, since a node with high betweenness centrality can control communication and the flow of information. Krackhardt also noted that a high standard betweenness centrality value can indicate that a node with high betweenness has access to a variety of non-redundant sources of information.
The relationship between degree centrality and closeness centrality was recently studied in [evans2022]. The authors found a non-linear relationship between degree centrality and closeness centrality. Degree, closeness, betweenness, eigenvector, and Katz centralities were mathematically examined with respect to whether these measures fulfill the authors’ three basic requirements for centrality measures in [landherr2010]. The correlation between community-aware centrality measures and classical centrality measures was studied in [rajeh2021].
Freeman discusses the notion that the center of a star-shaped network or the hub of a wheel-shaped network is structurally more central than any other node in any network of similar size [freeman1978]. However, this is more of a convincing argument than an exact mathematical definition. A comprehensive survey of centrality measures in complex networks gives a nice summary and overview of centrality measures [saxena2020centrality]. The survey highlights how degree centrality is a centrality measure based on local information, whereas closeness centrality and betweenness centrality make use of global knowledge of the network. Bovet and Makse give an overview of centrality measures based on whether they are based on shortest paths, walks, random walks, or non-backtracking walks [bovet2022centralities]. They note that degree centrality can actually be understood as the number of walks of length one starting from a node.
In a foundational work, Friedkin proposed a process model of social influence from which three centrality measures were defined [friedkin1991]. More recently, many new centrality measures have been proposed to identify influential nodes in complex networks [esfandiari2024collaborative, sadhu2024normalized, ghorbani2024opinion, bendahman2024unveiling, chiranjeevi2023icdc, qiu2021ranking].
5 Influence Spreading Model and New Centrality Metrics
In this section, we will introduce the influence spreading model, present a motivating example of how it works, and introduce new metrics defined consistently by the influence spreading model.
5.1 Influence Spreading Model
Our influence spreading model, previously published in [kuikkaSciRep], provides a detailed description of network structure and non-conserving probabilistic influence flow within that structure. The edges in the network can be directed or bidirectional, and both nodes and edges are assigned individual weights or probabilities for transmitting influence. The model supports both circular and recurrent spreading throughout the network. Additionally, a version of the model has been developed to implement only self-avoiding walks, meaning that each node can appear only once on a given walk. Furthermore, the model can account for breakthrough effects, enabling us to describe scenarios that lie between these two extreme cases.
The influence spreading model allows for two distinct types of spreading, namely complex contagion and simple contagion. Complex contagion allows unrestricted self-intersecting walks and cycles, with the only limitation for walks being a specified maximum walk length, and the target node cannot self-intersect. However, simple contagion does not allow for self-intersecting walks or cycles. In complex contagion, the focus is on the first arrival or the first influence on a target node. Before reaching the target node, the walks are limited by the specified maximum length. The calculation for each walk ends when it reaches the target node. In this article, we focus only on complex contagion since we wish to study centrality measures related to unrestricted paths, walks, and cycles and compare them to standard centrality measures.
The probability of spreading from node $s$ to node $t$ is calculated by considering all possible walks that are shorter than or equal to a specified maximum allowed walk length $L_{max}$. We assume that the spreading process starts with probability one from the initial node $s$. It is also important to emphasize that the influence spreading model is an analytical model and not a simulation model. Due to the analytical nature of the model, its probability outputs are exact within the limitations of floating-point number calculations.
The complete logic of calculating influence spreading probabilities was originally presented in [kuikka2022efficiency]. Depending on the context, the original article often used the term path interchangeably to mean either a path or a walk as defined in Section 1. In this article, we will aim to use more exact terminology that aligns with existing graph theory terminology.
The full influence spreading model specifies a time-dependent component in addition to edge and node probabilities [kuikka2022efficiency]. In the full model, this time-dependent probability component approaches one as time grows unbounded. In this article, we assume that the time component probability equals one. By setting the time component probability to one, we can have a clearer focus on how the influence spreading model’s centrality measures behave as the edge probability is varied.
In [kuikka2022efficiency], a detailed algorithmic description was given on how different walks and their probabilities are combined. Next, we will explain the logic of the walk combination process and the algorithm. Understanding how the influence spreading probability is calculated when walks are only restricted by the maximum allowed length of the walk is crucial to understanding the results that are presented in Section 6.
The probabilities of two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are combined by using their longest common prefix (LCP) walk $\mathcal{L}_{3}$ and the rules of probability theory. The reasoning behind this is to ensure that the results can be interpreted as valid probabilities. If $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ do not share any common walk, then we set $\mathbb{P}(\mathcal{L}_{3})=1$, where we denote probability with $\mathbb{P}$. The influence spreading model assumes that two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are conditionally independent given a common prefix walk $\mathcal{L}_{3}$ i.e. $\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2}\mid\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{1}\mid\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{2}\mid\mathcal{L}_{3})$. The formula for combining the probabilities of walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ given a common prefix walk $\mathcal{L}_{3}$ is derived as follows. We have
$\mathbb{P}(\mathcal{L}_{1}\cup\mathcal{L}_{2})$ $=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2})=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2}\cap\mathcal{L}_{3})$
$=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2}\mid\mathcal{L}_{3})$
$=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{1}\mid\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{2}\mid\mathcal{L}_{3})$
$=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{3})\frac{\mathbb{P}(\mathcal{L}_{1})}{\mathbb{P}(\mathcal{L}_{3})}\frac{\mathbb{P}(\mathcal{L}_{2})}{\mathbb{P}(\mathcal{L}_{3})}$
$=\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\frac{\mathbb{P}(\mathcal{L}_{1})\mathbb{P}(\mathcal{L}_{2})}{\mathbb{P}(\mathcal{L}_{3})},$ (4)
where the second equality follows from walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ both sharing the common walk $\mathcal{L}_{3}$, the third equality follows from the definition of conditional probability, the fourth equality follows from assumed conditional independence, and the fifth equality follows from $\mathbb{P}(\mathcal{L}_{i})=\mathbb{P}(\mathcal{L}_{i}\cap\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{i}\mid\mathcal{L}_{3})$ for $i=1,2$.
Equation (4) for combining two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ forms the basis of the influence spreading algorithm 1. Algorithm 1 and a more efficient algorithm for calculating influence spreading were originally presented in [kuikka2022efficiency]. For ease of understanding, we present the version which explicitly lists all possible allowed walks, although it is computationally less efficient. The computational limitations of algorithm 1 were discussed at length in [kuikka2022efficiency]. In practice, we have used the more efficient algorithm in our experiments. It is important to emphasize that both algorithm versions produce the exact same results.
Algorithm 1 Influence Spreading Algorithm
1:procedure Spread($s,t,G$)
[…]
3:  if $\mathcal{K}=\emptyset$ then return 0
[…]
5:  for $c=L_{max}\dots 0$ do
[…]
7:     continue
[…]
9:     $\mathbb{P}(\mathcal{L}_{1})\leftarrow\mathbb{P}(\mathcal{L}_{1})+\mathbb{P}(\mathcal{L}_{2})-\mathbb{P}(\mathcal{L}_{1})\mathbb{P}(\mathcal{L}_{2})/\mathbb{P}(\mathcal{L}_{3})$ $\triangleright$ Equation (4)
10:     $\text{Remove }\mathcal{L}_{2}\text{ from }\mathcal{K}$
11:  return $\mathbb{P}(\mathcal{L}_{1})$ $\triangleright$ The probability of the only walk left in $\mathcal{K}$
The influence spreading algorithm works by exploring all possible walks from source node $s$ to target node $t$ with at most $L_{max}$ edges. The possible walks are then sorted in increasing lexicographic order (shorter walks first; walks with the same length are sorted according to where they first differ). The walk combination process is done by first combining the walks with the longest common prefix (LCP). Then shorter LCPs are considered until all walks have been combined into one. Walks having an LCP of the same length can be combined in any order. In probabilistic terms, the actual combination of walk probabilities is calculated using equation (4).
5.2 Motivating Example of the Influence Spreading Model
We will now present an illustrative and motivating example of how the influence spreading algorithm works. This example will not only explain the influence spreading model, but it will also give insight into the results presented in Section 6.
Consider a simple four-node network presented in Figure 1, where edge weights denote influencing probabilities. While the network in Figure 1 is very simple, it shows some key features of the influence spreading model. Here we calculate the influence spreading probabilities for the maximum walk length of $L_{max}=5$. We consider two separate examples, first using node $1$ as the source and node $4$ as the target, and secondly using node $4$ as the source and node $1$ as the target. These calculations result in different end probabilities since there are more cycle possibilities when going from node $1$ to node $4$ than vice versa.
31240.50.50.50.5
Figure 1: A simple network where the edge weights are probabilities.
When node $1$ is the source and node $4$ is the target, we have in total ten different walks that have length 5 or less. These ten walks are illustrated in Figure 2. In contrast, when node $4$ is the source and node $1$ is the target, then we have in total only six different walks that have length 5 or less. These six walks are illustrated in Figure 3.
3141.2.
31241.2.3.
31241.2.3.4.
3141.2.3.4.
31241.2.3.4.
31241.2.3.4.5.
31241.2.3.4.5.
31241.2.3.4.5.
31241.2.3.4.5.
31241.2.3.4.5.
Figure 2: All ten walks from node 1 to node 4 with walk length equal or less than 55. The numbers denote order of traversal along the edges of the network.
3141.2.
31241.2.3.
31241.2.3.4.
3141.2.3.4.
31241.2.3.4.5.
31241.2.3.4.5.
Figure 3: All six walks from node 4 to node 1 with walk length equal to or less than 55. The numbers denote the order of traversal along the edges of the network.
When the influence spreading algorithm 1 is performed with node 4 as the source and node 1 as the target, we find six walks with the walk length shorter than or equal to $5$. The walks are then sorted in increasing lexicographic order as follows:
  1. 1.
    $4\rightarrow 3\rightarrow 1$
  2. 2.
    $4\rightarrow 3\rightarrow 2\rightarrow 1$
  3. 3.
    $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$
  4. 4.
    $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 1$
  5. 5.
    $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$
  6. 6.
    $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$.
The longest common prefix (LCP) walks in this lexicographic order are walks including $4\rightarrow 3\rightarrow 2\rightarrow 3$ and $4\rightarrow 3\rightarrow 4\rightarrow 3$. We can process walks with the same LCP walk lengths in any order. An example calculation of the combination of walk probabilities is combining the third $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$ and fifth $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$ walks. The third walk has probability $0.5^{4}$, the fifth walk has probability $0.5^{5}$, and their LCP walk has probability $0.5^{3}$. Therefore, the combined probability is $0.5^{4}+0.5^{5}-0.5^{4}\cdot 0.5^{5}/0.5^{3}=0.078125$. When the algorithm 1 is executed until completion, we get the probability of approximately $0.358$. This means that node 4 influences node 1 with an approximate probability of $0.358$. When the influence spreading algorithm 1 is performed with node 1 as the source and node 4 as the target, we get the result that node 1 influences node 4 with approximately $0.472$ probability. Hence, we see that node 1 has more influence on node 4 than vice versa. This result is explained by the different number of walk possibilities.
Figures 2 and 3 underscore a crucial aspect of the influence spreading model: Calculations are conducted only until the target node is reached in each walk. This feature exists because the influence spreading model focuses on the first arrival/influence. This feature results in a different number of possible walks when going from node 1 to node 4 and vice versa. There are four fewer paths from node 4 to 1 than there are from node 1 to 4. The position of node 1 gives it more possibilities for cycles before reaching the target node 4. In contrast, the position of node 4 allows for fewer cycle possibilities in the network before reaching target node 1. We also see that node 4 is on the periphery of this small network, while node 1 is more central. This is also an important observation when we discuss the results of other experiments in Section 6, where we will also refer to a more formalized definition of the periphery.
5.3 New Centrality Metrics
The output of the influence spreading model is represented as a matrix, known as the influence spreading matrix. The elements of this matrix describe the directed probabilities of influence spreading from a source node $i$ to a target node $j$ for all pairs of $N\times N$ nodes within a network consisting of $N$ nodes. The elements of the matrix reflect the probabilities of directed spreading or connectivity between nodes in the network $V$, and it is used to analyze the structure and flow processes of the network. Importantly, models other than our influence spreading model can also create this probability matrix. Using the influence spreading matrix $M$, we define our new out-centrality $C_{out}$ and in-centrality $C_{in}$ as follows
$C_{out}(i)=\frac{1}{N-1}\sum_{j\in V\setminus\{i\}}M(i,j),\quad i=1,...N$ (5)
and
$C_{in}(j)=\frac{1}{N-1}\sum_{i\in V\setminus\{j\}}M(i,j),\quad j=1,...,N.$ (6)
We assume that the spreading process begins with the probability of one from the initial node. When spreading occurs from branching paths or walks, it can happen in multiple directions, indicating the likelihood of first arrival at the end nodes. By ignoring the diagonal elements of the influence spreading matrix (a node’s influence on itself is always one), we can define the out-centrality and in-centrality values. To normalize these values to one (or express them as percentages), we use the factor 1/(N1)1/(N-1) on the right side of equations (5) and (6). Out-centrality measures how much a node influences other nodes on average, and in-centrality measures how much a node is influenced by other nodes on average.
The influence spreading model also defines its own betweenness centrality measure. We refer to this betweenness as ISM betweenness centrality or ISM betweenness, which is defined with the help of the concept of cohesion. The cohesion $\mathcal{C}$ of the network $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges in the network, is defined as C = ∑_s, e ∈V s ≠e M(s, e). The cohesion $\mathcal{C}_{v}$ of the network without node $v$ and its removed incident edges is defined as C_v = ∑_s, e ∈V ∖{ v } s ≠e M^*(s, e), where $M^{*}$ is the influence spreading matrix calculated from the network without node $v$ and its incident edges. With these definitions, we define ISM betweenness centrality $b_{ISM}(v)$ of a node $v$ as
$b_{ISM}(v)=\frac{\mathcal{C}-\mathcal{C}_{v}}{\mathcal{C}}.$ (7)
It is important to emphasize the way the ISM betweenness $b_{ISM}$ is calculated. When the ISM betweenness value of node $v$ is calculated, we remove the node and its incident edges from the network, and then the logic of the influence spreading algorithm 1 is performed on this altered network. It is important to note that ISM betweenness measures the relative importance of node $v$ by examining how much the cohesion $\mathcal{C}$ of the network changes when node $v$ is removed. If node $v$ acts as a particularly important bridge in the network, its removal can even lead to the network being disconnected into two separate components.
6 Results
In this section, we will compare the metrics defined by the influence spreading model to standard centrality metrics. Furthermore, we will further examine the behavior of the centrality metrics defined by the influence spreading model.
The influence spreading model defines its own betweenness centrality (7) and also new centrality metrics out-centrality (5) and in-centrality (6). The compared standard centrality metrics are betweenness centrality (3), degree centrality (1) and closeness centrality (2).
The influence spreading model uses probabilities as edge weights. In our experiments, the edge probability is varied between experiments, but all edges of the network always have the same probability. However, the standard betweenness centrality and the closeness centrality require a distance to be defined between the nodes. We interpret the shortest path between the nodes as the path with the fewest edges (as if the distance between the nodes were one). If edge probabilities are not the same for all edge weights, then a systematic way of mapping probabilities to distances is needed. Appendix A explores one possible way of mapping edge probabilities to edge distances. However, in our experiments, the shortest paths are calculated with the assumption that all edge distances are equal. Therefore, calculating the shortest path between nodes reduces to finding the path with the fewest edges.
We frequently characterize nodes as peripheral or central. Borgatti and Everett proposed formalizations for the concept of core/peripheral structure in networks [borgatti2000]. One of the formalizations they proposed was a discrete model consisting of a core and a periphery. The core was defined as a cohesive subgraph in which nodes are connected to each other in a maximal sense, and the periphery consists of nodes that are loosely connected to the core but do not share the maximal cohesion with the core. Yanchenko and Sengupta explain the core as consisting of nodes that are densely connected to each other, and some of them also have connections to peripheral nodes, whereas peripheral nodes are only sparsely connected to each other [yanchenko2023]. We use the term peripheral nodes in the sense and meaning given by Yanchenko and Sengupta, and central nodes are interpreted as belonging to the core.
In Sections 6.1 and 6.2, we will have a detailed individual node-level examination of the behavior of the influence spreading model and its centrality metrics on two small networks. Using small networks allows us to examine and explain the behavior of the influence spreading model and its centrality metrics with node-level detail. In Section 6.3, we examine the collective network-level behavior of the influence spreading model and its centrality metrics on larger $1000$ node random graphs/networks. Finally, in Section 6.4, we show a few counterexamples where some of the phenomena we have found do not appear. All experiments in Sections 6.1, 6.2, and 6.3 use the maximum allowed walk length $L_{max}=20$. In Section 7, the effect of using shorter or longer $L_{max}$ is discussed.
6.1 Detailed Node-Level Analysis of the Influence Spreading Model on the Krackhardt Kite Network
We will examine the behavior of the influence spreading model in the Krackhardt kite network [krackhardt1990]. The Krackhardt kite shown in Figure 4 is a useful example to compare the influence spreading model against standard centrality metrics, as it reveals the differences of the influence spreading model’s centrality measures compared to standard centrality measures. The Krackhardt kite network is well known for producing a different highest ranked node for each of the calculations of standard betweenness (node 8), closeness (nodes 6 and 7), and degree centrality (node 4) [krackhardt1990].
13425678910
Figure 4: Krackhardt kite. This network was originally presented in [krackhardt1990].
In Figure 5, we see the values and node rankings for the betweenness centrality measure as defined by the influence spreading model. The rankings of ISM betweenness are plotted side by side with the standard betweenness centrality rankings. The ISM betweenness centrality value is dependent on the edge probability. Figure 5 also shows how the rankings between nodes change as the edge probability increases. Standard betweenness gives different rankings compared to ISM betweenness for all edge probabilities. The greatest difference is how nodes 8, 9, and 4 behave. Their rankings change quite markedly when the edge probability is increased.
Figure 5: Influence spreading model betweenness values and rankings alongside standard betweenness centrality rankings for the Krackhardt kite network displayed in Figure 4. Node labels are marked alongside the plotted points.
It is also important to note how ISM betweenness gives non-zero values for nodes 3, 5, and 10, whereas standard betweenness assigns them the value zero since they do not appear in any shortest path between nodes. However, by allowing any walks including cycles, the influence spreading model recognizes that these nodes could have more importance than mere shortest path calculations could imply.
When edge probabilities are one, the ISM betweenness values converge into three separate values. This phenomenon is explained by how the ISM betweenness is calculated. Node 8 attains the highest ISM betweenness value for large edge probability values. When node 8 and its incident edges are removed from the network, the network is divided into two separate components: one containing two nodes and the other containing seven nodes. This results in a large difference in the cohesion values $\mathcal{C}$ and $\mathcal{C}_{8}$, and therefore the ISM betweenness value is high. A similar phenomenon occurs when node 9 and its incident edges are removed from the network. Again, the network is divided into two components, one containing only one node and the second containing eight nodes. For all other nodes, removing them from the network does not create two separate components. Instead, the network stays connected, which allows all influences to be maximally transferred between all nodes. Therefore, for nodes 1, 2, 3, 4, 5, 6, 7, and 10, their ISM betweenness values converge to the same value.
In Figure 6, the node rankings given by out-centrality and in-centrality are compared to the rankings given by standard closeness and degree centrality. Since out-centrality measures a node’s average influence on other nodes and in-centrality measures the other nodes’ average influence on a node, these centrality metrics measure different things. Additionally, standard closeness centrality measures how close nodes are to each other and degree centrality measures the number of neighbors a node has. Therefore, all of these centrality metrics measure different things, but despite this, we notice similarities between these metrics. Out-centrality rankings remain unchanged as the edge probability increases. We see that in rankings, there is great similarity between out-centrality and standard closeness centrality, and out-centrality and degree centrality.
Figure 6: Influence spreading model out-centrality and in-centrality rankings compared to standard closeness and degree centrality rankings for the Krackhardt kite network displayed in Figure 4. Node labels are marked alongside the plotted points.
In Figure 6, the in-centrality ranking is initially exactly the same as the out-centrality ranking. Thus, for low edge probability values, there is a great similarity to closeness centrality and degree centrality. However, as the edge probability value is increased, the in-centrality rankings start to reverse. From the edge probability value $0.4$ onward, the rankings have completely reversed. Consequently, for edge probabilities that are equal to or greater than $0.4$, there is a reverse order relationship between in-centrality and closeness centrality and between in-centrality and degree centrality.
In Figure 7, we see the values for the out-centrality and in-centrality measures defined by the influence spreading model. From small to large edge probabilities, out-centrality values have much larger dispersion than in-centrality values. In Figure 6, we saw how the in-centrality rankings initially developed in a similar fashion to out-centrality rankings, but for higher edge probabilities, the rankings were completely reversed. Figure 7 tells the same story with one important difference as it highlights how the value differences in in-centrality for the Krackhardt kite network are small. However, the small differences in the in-centrality values are still significant since they are explained by the network structure.
Additionally, a notable feature of the influence spreading model is also shown in Figure 7. When edge probabilities are $1.0$ for all edges, all of the influence is coming through, and every node attains a trivial $1.0$ value for their out- and in-centrality measures. When edge probabilities are almost zero for all edges, almost none of the influence is coming through, and every node achieves a value close to zero for their out- and in-centrality measures.
Figure 7: Influence spreading model out-centrality and in-centrality values for the Krackhardt kite network displayed in Figure 4. The lowest probability is $0.01$ and thereafter probabilities increase from $0.1$ to $1.0$ in $0.1$ increments.
In Figures 6 and 7, we see that peripheral nodes attain larger in-centrality values for high edge probabilities than more central nodes. The reason for this phenomenon is found in the example of Section 5.2. When a more central node is exerting influence towards a peripheral node, it has more cycle possibilities than a peripheral node due to its position in the network. In contrast, when a more peripheral node is exerting influence towards a central node, it has fewer cycle possibilities than a central node due to its position in the network. It is also important to emphasize that if a target node has a high degree, then all of the edges leading to the target node are only allowed as the last edge on a walk. The influence spreading model focuses on the first arrival (first influence) and therefore peripheral nodes are influenced more than central nodes when edge probabilities are high, since high edge probabilities allow the influence to reach a target node even for very long paths.
Figures 6 and 7 demonstrate that the out-centrality ranking between nodes remains the same, although their values change as the edge probability value increases. This result can be interpreted as the edge probability acting as an amplifier to a node’s position in the network. Every node can exert more influence, but the more central nodes still have more cycle possibilities than the more peripheral nodes. Therefore, the node rankings remain completely unchanged.
6.2 Detailed Node-Level Analysis of the Influence Spreading Model on Another Small Network
In Figure 8, a small 10-node network is presented. The same network was used in [landherr2010] to compare standard centrality measures with each other. This small network has an almost chain-like structure, where nodes exist mainly on a chain with some exceptions. Applying the influence spreading model on this small network results in unique values for all the centrality measures defined by the influence spreading model when the common edge probabilities are not equal to zero or one.
12345678109
Figure 8: A small network with 10 nodes. This network was originally used in [landherr2010].
In Figure 9, ISM betweenness values and rankings are plotted, and the rankings are compared to standard betweenness values. It is notable that the rankings shown in the figure are similar to each other. However, there are some notable differences. The standard betweenness centrality assigns value 0 to nodes 1, 8, and 3. In contrast, ISM betweenness for node 3 is higher than for nodes 1 and 8 for all edge probabilities. Since the influence spreading model also accounts for walks and paths that are not the shortest path, this distinction makes sense. Node 3 is indeed between other nodes on certain walks and paths that are not the shortest paths. Therefore, the influence spreading model finds information that using only the shortest paths cannot.
Figure 9: Influence spreading model betweenness values and rankings alongside standard betweenness centrality rankings for the small network displayed in Figure 8. Node labels are marked above the lines.
In Figure 9, we see that for the edge probability value $1$, the ISM betweenness values show four separate and distinct values, and in some cases, multiple nodes have an ISM betweenness value that converges to the same value. Nodes 5 and 6 attain the highest ISM betweenness values for high edge probabilities and converge at the edge probability value $1$. When either node 5 or 6 and its incident edges are removed from the network, the network is separated into two components, one containing four nodes and the other containing five nodes. Removing node 4 and its incident edges results in the first component containing three nodes and the second component containing six nodes. Removing node 2 or 7 results in the first component that contains one node and the second component that contains eight nodes. Hence, nodes 2 and 7 have converging ISM betweenness values. Nodes 1, 3, 8, 9, and 10 have converging ISM betweenness values as well, since the removal of any of them and their incident edges results in the network still staying connected.
In Figure 9, there are also other notable phenomena. Node 5 is ranked as the second node when standard betweenness is used. Using ISM betweenness, node 5 attains this same rank only for higher edge probabilities. This can be interpreted such that for lower edge probabilities, the influence from and to surrounding node clusters on both sides (cluster of nodes 2, 3, and 4 and cluster of nodes 6, 7, 9, and 10) does not come through very well. However, for higher edge probabilities, the influence coming from and to node clusters on each side of it comes through with more intensity. The other notable phenomenon is the behavior of the nodes (3, 9, and 10) that are not positioned on the spine of the chain-like network structure. Initially, nodes 9 and 10 have lower ISM betweenness values than node 3, but as the edge probability value increases, this situation is eventually reversed for the edge probability value $0.9$.
In Figure 10, the out-centrality and in-centrality rankings are compared to the standard closeness centrality and degree centrality rankings. Figure 10 shows that there is a similarity in the rankings between out-centrality and standard closeness centrality. Initially, in-centrality also has similarity to standard closeness centrality since its values are initially the same as out-centrality. Many nodes of the network have the same degree and therefore some rankings are shared between nodes. Shared rankings complicate comparisons, but we still see how the highest degree nodes 2, 4, 6, and 7 also attain the highest four ranks of out-centrality values for all edge probabilities. Similarly, nodes 3, 5, 9, and 10 are jointly ranked fifth for degree centrality, and they attain out-centrality ranks ranging from fifth to eighth for all edge probabilities. The peripheral nodes 1 and 8 are at the bottom of the rankings for both out-centrality and degree centrality.
Figure 10: Influence spreading model out-centrality and in-centrality rankings compared to standard closeness and degree centrality rankings for the small network displayed in Figure 8. Node labels are marked above the lines.
In Figure 10, we see again the reversing in-centrality ranking behavior first seen for the Krackhardt kite network in Figure 6. Although out-centrality and in-centrality initially share the same rankings, as edge weight increases, the relationship begins to reverse. With the highest edge probabilities, the in-centrality rankings have fully reversed compared to out-centrality rankings.
Figure 10 shows us that initially, nodes on the periphery of the network (nodes 1 and 8) are influenced the least amount by other nodes simply because the influence of other nodes hardly reaches them due to low edge probabilities. When edge probabilities reach a sufficiently high level, the peripheral nodes become the most influenced nodes in the network.
On the other hand, Figure 10 shows that out-centrality rankings stay relatively stable for all edge probabilities. What this means is that the most influential nodes remain most influential regardless of the edge probability value. However, there is a notable phenomenon in the out-centrality rankings. Nodes 6, 7, and 8 switch places with nodes 4, 2, and 1, respectively, as the edge probability value increases. Nodes 4 and 6, nodes 2 and 7, and nodes 1 and 8 have similar positions in the network chain. Initially, at low edge probabilities, node 6 ranks higher than node 4, node 7 ranks higher than node 2, and node 8 ranks higher than node 1. One possible explanation for this is that nodes 6, 7, and 8 have greater closeness centrality compared to their counterparts—nodes 4, 2, and 1, respectively. This higher closeness centrality allows nodes to exert more influence, even at lower edge probabilities. Beginning at the edge probability of $0.5$, nodes 4, 2, and 1 start to surpass their counterpart nodes 6, 7, and 8, respectively.
Figure 10 also shows that for low edge probabilities, nodes on the periphery (nodes 1 and 8) receive relatively little influence from other nodes since their in-centrality rankings are low. We can infer that for low edge probabilities, the most influential nodes mainly influence nodes close to them, and their influence does not reach the periphery.
In Figure 11, we see the out-centrality and in-centrality values for the small network. It is notable that, unlike the Krackhardt kite network, out-centrality values have less dispersion. This reduced dispersion between values is due to the chain-like structure of this small network, which does not allow nodes to have wildly different positions in the network.
Figure 11: Influence spreading model out-centrality and in-centrality values for the small network displayed in Figure 8. The lowest probability is $0.01$ and thereafter probabilities increase from $0.1$ to $1.0$ in $0.1$ increments.
6.3 Network-Level Analysis of the Influence Spreading Model in Random Networks
In this section, we examine the behavior of the influence spreading model on larger randomly generated networks and compare the results to standard centrality metrics. We use scatter plots and Pearson correlation coefficients in our comparisons. The sample Pearson correlation coefficient $r_{xy}$ for a sequence of observation pairs $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ is defined as
$r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\left(\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}\right)\left(\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}\right)}},$
where $\bar{x}=(1/n)\sum_{i=1}^{n}x_{i}$ and $\bar{y}=(1/n)\sum_{i=1}^{n}y_{i}$ [kaptein2022].
We have used stochastic network models to generate the random networks used in this section. The stochastic network models used are Erdős-Rényi random graph model [erdos1959], Watts-Strogatz small world model [watts1998] and Barabási-Albert preferential attachment model [barabasi1999]. Erdős-Rényi model comes from the mathematical tradition of random graph theory, whereas Watts-Strogatz and Barabási-Albert models are innovations of network science. These three random network models each capture a different type of network with their characteristic properties and provide an aggregate representation of many individual networks that share these properties. The characteristic properties of the random networks produced by these stochastic network models are very well known, and they have been extensively studied in the literature. Therefore, these three random network models provide varied, yet controlled conditions to assess the influence spreading model and its centrality metrics compared to standard centrality metrics.
We also note that we performed the same or similar experiments on the well-known Les Misérables network [knuth1993] and on another random network generated by using the extension of the Barabási-Albert model proposed by Holme and Kim [holme2002]. However, the Les Misérables network data set is quite small with 77 nodes and 254 edges, while similar phenomena appeared in the results as in our chosen three random networks. Therefore, the inclusion of the Les Misérables network was considered redundant. Furthermore, the random network produced by the method proposed by Holme and Kim displayed results similar to Barabási-Albert random network. This is not entirely surprising, since the essential difference between these random network models is that the method of Holme and Kim can create high clustering, whereas the Barabási-Albert model cannot.
The model widely known as the Erdős-Rényi model (or less widely known as the Erdős-Rényi-Gilbert model) has two formulations [erdos1959, gilbert1959]. In the first formulation, the number of nodes and edges is fixed, and then a graph is randomly selected out of all possible graphs. In the second formulation, the number of nodes is fixed, and edges are created between all pairs of nodes independently of all other edges with probability $p$. In this study, we use the second formulation and denote a random Erdős-Rényi network with $\text{ER}(n,p)$, where $n$ is the number of nodes, and $p$ is the independent edge creation probability.
The Erdős-Rényi model has been thoroughly studied and its properties are well known [newman2003]. The Erdős-Rényi model creates networks with low clustering, and its degree distribution has been shown to be the Poisson distribution.
The Watts-Strogatz model uses a regular ring lattice of (fixed number of) $n$ nodes and $k$ neighbors for each node as the starting point, after which edges are rewired with probability $p$ [watts1998]. Thus, with probability $p$, a node is connected to a uniformly randomly selected non-neighbor (duplicate edges are not allowed) and with probability 1p1-p the edge remains unchanged. With this method of network construction, Watts and Strogatz aimed to develop a mixture of order and randomness to create more realistic networks. The parameter $p$ controls the amount of randomness. The Watts-Strogatz model has the ability to create high clustering and short average path lengths between nodes. Due to the short average path lengths, it is also called the small-world model. The exact degree distribution of the Watts-Strogatz model is derived in [albert2002]. When $p=0$, the degree distribution is constant $k$. When $p$ grows, the degree distribution begins to resemble the shape of the degree distribution of the Erdős-Rényi random graph [albert2002] i.e. the degree distribution becomes approximately Poisson distributed. We denote a network produced by the Watts-Strogatz model by $\text{WS}(n,k,p)$.
The Barabási-Albert model addresses some of the limitations of Erdős-Rényi model and Watts-Strogatz model [barabasi1999]. First, the Barabási-Albert model does not assume a fixed number of nodes, and instead, the number of nodes is allowed to grow. Second, Barabási-Albert model does not create edges between nodes with equal probability since a new edge is more likely to be connected to a node that already has a high degree. This is called preferential attachment, where high-degree nodes are more likely to receive more neighbors when the network size grows. The Barabási-Albert model begins from an initial graph and then continues adding nodes using preferential attachment until the desired network size has been reached. We denote the Barabási-Albert network with $\text{BA}(n,m)$, where $n$ is the desired number of nodes, and $m$ is the number of edges to be attached from a new node to existing nodes. The degree distribution of the Barabási-Albert model is scale-free.
Many real-world networks display high clustering and non-Poissonian degree distributions [newman2003]. Scale-free (i.e. power-law) degree distributions have been proposed to be a common phenomenon for many real networks [albert2002]. However, the prevalence of scale-free degree distributions in real networks has also been recently disputed in [broido2019], where it is actually found that scale-free degree distributions are rare and the log-normal distribution provides an adequate modeling fit instead of a power-law distribution. A short overview on the prevalence of scale-free degree distributions in real networks can be found in [holme2019]. In light of this differing research on what degree distributions characterize real networks, we argue that testing the behavior of the influence spreading model on random networks that display either Poissonian or scale-free degree distributions is especially justified.
We created three separate random networks with $1000$ nodes each using the aforementioned Erdős-Rényi, Watts-Strogatz and Barabási-Albert models. All random networks were created using the Python package NetworkX  [SciPyProceedings_11]. The parameters were chosen such that each random network would have roughly the same number of edges. The random network $\text{ER}(1000,0.01)$ contains $1000$ nodes and $4910$ edges. The random network $\text{WS}(1000,10,0.5)$ has $1000$ nodes and $5000$ edges. The random network $\text{BA}(1000,5)$ contains $1000$ nodes and $4975$ edges. All three random networks are connected. The networks were created using seed $1729$.
In Figure 12, we have compared ISM betweenness centrality to standard betweenness centrality for three random networks produced by the Erdős-Rényi, Watts-Strogatz, and Barabási-Albert models. Figure 12 reveals how, for lower edge probabilities, there is quite a strong positive correlation between the ISM betweenness and the standard betweenness. As the edge probability value increases, this correlation becomes weaker and weaker. Eventually, for the edge probability value $0.9$, we see that multiple levels are formed for the ISM betweenness values. This phenomenon is likely caused by differences in how the network breaks down into components when a node is removed.
Figure 12: Influence spreading model (with maximum walk length $L_{max}=20$) betweenness centrality compared to standard betweenness centrality for edge probabilities $0.1,0.3,0.5,0.7$ and $0.9$ on different random networks.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.93 0.94 0.80 0.63 0.45 0.30 0.18 0.11 0.06
$\text{WS}(1000,10,0.5)$ 0.89 0.89 0.79 0.66 0.53 0.38 0.25 0.16 0.10
$\text{BA}(1000,5)$ 0.96 0.88 0.86 0.85 0.84 0.82 0.73 0.42 0.17
Table 1: Pearson correlation coefficient between ISM betweenness centrality and standard betweenness centrality.
In Table 1, we see that the $\text{BA}(1000,5)$ network shows quite strong to moderate positive correlation values for higher edge probabilities, whereas the correlation has significantly weakened for the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks. Eventually, positive correlation becomes weak for the $\text{BA}(1000,5)$ network as well.
Interestingly, the extension of the Barabási-Albert model proposed by Holme and Kim produced a random network that maintained a strong correlation between ISM betweenness and standard betweenness even for the highest edge probability value $0.9$. Stronger positive correlation between ISM betweenness and standard betweenness for the edge probabilities $0.7,0.8$, and $0.9$ was the only notable difference of the Holme-Kim network compared to the $\text{BA}(1000,5)$ network in our experiments.
In Figure 13, out-centrality is compared to degree centrality. The Barabási-Albert model creates a heavy-tailed degree distribution (scale-free distributions are heavy-tailed), resulting in hubs. A hub is a node with a great number of neighbors compared to the average degree. In this particular 1000-node Barabási-Albert network, there is a node with more than 100 neighbors. This and other hubs cause a non-linear relationship between out-centrality and degree centrality values even for the edge probability value $0.1$.
Figure 13: Influence spreading model (with maximum walk length $L_{max}=20$) out-centrality compared to degree centrality for edge probabilities $0.1,0.3,0.5,0.7$, and $0.9$ on random networks.
Another notable phenomenon in Figure 13 is how the scale of out-centrality values changes for different networks as the probability of the edges increases. Both networks $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ display very small out-centrality values for the edge probability value $0.1$. However, for higher edge probabilities, some nodes attain out-centrality values that are close to one (maximum value for out-centrality), and as the edge probability increases, the interval of out-centrality values begins to narrow as minimum values approach one. For the $\text{BA}(1000,5)$ network, the pattern is similar with one notable exception. For the edge probability value $0.1$, the hub nodes of this network already achieve values that are close to the maximum out-centrality value.
In Table 2, the Pearson correlation coefficient between out-centrality and degree centrality is shown for all edge probabilities. For the edge probability value $0.1$, the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display a quite strong positive correlation, which then becomes weaker as the edge probability increases. For the $\text{BA}(1000,5)$ network, there is also a positive correlation. However, the non-linear relationship makes the correlation weaker.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.95 0.93 0.82 0.69 0.55 0.42 0.31 0.22 0.16
$\text{WS}(1000,10,0.5)$ 0.93 0.96 0.90 0.82 0.70 0.57 0.43 0.31 0.23
$\text{BA}(1000,5)$ 0.87 0.68 0.57 0.49 0.41 0.29 0.15 0.06 0.03
Table 2: Pearson correlation coefficient between out-centrality and degree centrality.
Figure 13 and Table 2 show that for low edge probabilities, the degree is a good predictor for out-centrality values when the degree distribution is Poissonian. When the degree distribution is scale-free, a linear relationship does not fully capture the dynamics between out-centrality and degree centrality. For higher edge probabilities, all the random networks begin to display a non-linear relationship between out-centrality and degree centrality. $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks maintain a stronger positive correlation with degree centrality for much higher edge probabilities than $\text{BA}(1000,5)$ network. The difference again comes from the light-tailed Poisson degree distribution that is shared by the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks, whereas $\text{BA}(1000,5)$ network has a heavy-tailed degree distribution.
In Figure 14, out-centrality is compared to the standard closeness centrality for $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ random networks. Again, it is important to notice the changing scale of out-centrality values as the edge probability increases. Out-centrality and standard closeness centrality display a similar relationship to what was seen between out-centrality and degree centrality.
Figure 14: Influence spreading model (with maximum walk length $L_{max}=20$) out-centrality compared to standard closeness centrality for edge probabilities $0.1,0.3,0.5,0.7$, and $0.9$ on random networks.
In Table 3, we see a positive correlation for low edge probabilities, which becomes weaker as the edge probability increases, and eventually the correlation is close to zero. As edge probability increases, $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks maintain a stronger positive correlation than $\text{BA}(1000,5)$ network. We see that for low edge probabilities, closeness centrality is a good predictor for out-centrality values. We can see that for low edge probabilities, the close and immediate neighborhood of a node determines its ability to influence other nodes (out-centrality), since the low influence probability decays significantly over longer walk distances.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.97 0.96 0.87 0.76 0.63 0.50 0.39 0.29 0.21
$\text{WS}(1000,10,0.5)$ 0.89 0.87 0.82 0.74 0.65 0.54 0.42 0.32 0.25
$\text{BA}(1000,5)$ 0.92 0.76 0.65 0.56 0.47 0.33 0.18 0.07 0.03
Table 3: Pearson correlation coefficient between out-centrality and standard closeness centrality.
In Figure 15, the relationship between in-centrality and out-centrality is studied for the $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ networks. It is important to note that both the x-axis and y-axis scale changes in every individual plot. We see the same behavior for out-centrality and in-centrality values that was seen for the Krackhardt kite and small network examples in Sections 6.1 and 6.2, respectively. Out-centrality values have much larger dispersion, whereas there is a small variation between in-centrality values.
Figure 15: Influence spreading model (with maximum path length $L=20$) in-centrality compared to out-centrality for edge probabilities $0.1,0.3,0.5,0.7$, and $0.9$ on random networks.
In Figure 15 and in Table 4, both $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display an initially strong positive correlation, which eventually reverses into a strong negative correlation as the edge probability increases. However, for the $\text{BA}(1000,5)$ network, the reversal has already begun for the edge probability value $0.1$ since part of the scatter plot shows a positive correlation, while the other part shows a negative correlation. As the edge probability increases, the correlation between in-centrality and out-centrality becomes strongly negative. We are again essentially seeing the inverting ranking behavior of out-centrality and in-centrality as the edge probability increases.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.99 -0.98 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
$\text{WS}(1000,10,0.5)$ 0.99 -0.99 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
$\text{BA}(1000,5)$ 0.53 -0.97 -0.99 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
Table 4: Pearson correlation coefficient between in-centrality and out-centrality.
6.4 Special Cases
In Figure 16, a tetragon network is shown on the left and the same network with one peripheral node added is shown on the right. For the regular tetragon network, given an identical edge probability, each node attains the same ISM betweenness centrality value, the same out-centrality value, and the same in-centrality value. When a peripheral node is added, the results change in a way that node 4 becomes the highest ranked in ISM betweenness value, whereas node 5 behaves like the peripheral nodes we saw in our earlier two examples. Nodes 1 and 3 still attain identical values to each other, although their numerical values change.
4132
41325
Figure 16: A tetragon network on the left. When a peripheral node is added on the right, the results of the influence spreading model change dramatically.
The tetragon network shows that not all networks produce the inverting ranking behavior between out-centrality and in-centrality as was seen for the Krackhardt kite network in Figure 4, for the small node network in Figure 8 and for the larger random networks. It also further reinforces that when the network contains symmetry, the results of the influence spreading model are also symmetric, as long as the edge probabilities are the same for all edges.
Another case, when network-wide inverting ranking behavior between out-centrality and in-centrality does not happen, is when nodes are completely disconnected from the rest of the network. If a node is disconnected from the rest of the network i.e. it has degree zero, then its in-centrality and out-centrality equal zero for all edge probabilities. The rest of the network may exhibit a reversing behavior between in-centrality and out-centrality; however, nodes without neighbors disrupt the change in correlation from positive to negative, as their values remain static at zero.
A particular Dutch student network is an example of this phenomenon, where the correlation between in-centrality and out-centrality remains positive as the edge probability increases. The Dutch student network consists of 32 students, where their friendship was originally mapped and studied in [van1999], and the network is visualized in Figure 17. This student network contains three students (nodes) whose degree is zero. In actuality, the network has four components: one very large component and the three separated nodes, each forming its own component.
2621147162102231817242331525321911276192829301320451218
Figure 17: A Dutch student network with three nodes (5, 12, and 18) completely disconnected from all other nodes. Based on data collected by Van de Bunt [van1999]
In Figure 18, we see that although the rest of the nodes in the network display the reversing behavior between in-centrality and out-centrality as the edge probability increases, the three zero-degree nodes maintain their position as last in the rankings. If the nodes with degree zero were removed, we would observe the same inverted ranking behavior for the whole network that was seen before.
Figure 18: Out-centrality and in-centrality rankings for the Dutch student network shown in Figure 17. Node labels are plotted alongside the lines. Notably, multiple nodes share the same rank.
In Figure 19, the in-centrality and out-centrality values for this student network are plotted in a scatter plot. We can see that the positive correlation persists because there are three zero-degree nodes with both in-centrality and out-centrality values of zero.
Figure 19: In-centrality and out-centrality values plotted for increasing edge probabilities for the Dutch student network shown in Figure 17
In Table 5, we observe a positive correlation between in-centrality and out-centrality values across all edge probabilities. For edge probabilities in the middle range, the correlation becomes weaker but remains positive. However, for the highest edge probability, the correlation strengthens significantly, becoming strongly positive again. If we remove the three zero-degree outlier nodes, the correlation shifts from positive to negative, as illustrated by the ranking changes in Figure 19.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Students 1.00 0.98 0.63 0.48 0.60 0.75 0.87 0.95 0.99
Table 5: Pearson correlation coefficient between in-centrality and out-centrality for the Dutch student network shown in Figure 17.
7 Discussion
The influence spreading model serves as a versatile framework to investigate properties of various complex networks as it also allows for directed unidirectional or bidirectional edges and any edge probabilities to be used. In this article, however, we have studied undirected networks, where all edge probabilities are assumed to be equal. This setting has given us a controlled environment, where the effect of the edge probability can be captured more easily. We have also developed an understanding of how the influence spreading model behaves when the edge probability is varied.
All experiments performed in Sections 6.1, 6.2, and 6.3 were performed using the maximum allowed walk length $L_{max}=20$. We also performed all these experiments for shorter and longer maximum allowed walk lengths. The effect of the maximum length of allowed walk is that the reversing ranking behavior in the in-centrality occurs earlier if walks are allowed to be longer than the length of $20$, and reversing occurs later if walks are only allowed to be shorter than the length of $20$. In general, a longer $L_{max}$ increases the probability of influencing others, as it allows more walks within the network. In [kuikka2022efficiency], the effect of varying the maximum allowed walk length was discussed in detail (the article used the term path instead of walk while retaining the meaning of walk as defined in this article).
In this study, we have only presented the complex contagion version of the influence spreading model. For random networks produced by the Erdős-Rényi, Watts-Strogatz, and Barabási-Albert models, the complex contagion process results in in-centrality and out-centrality values being positively correlated for low edge probabilities. As the edge probability is increased, eventually this relationship between in-centrality and out-centrality completely reverses, where there exists a negative correlation between these two measures. It is important to emphasize that simple contagion cannot create this phenomenon since cycles are not allowed, and therefore there are the same number of paths from a source node to a target node as there are from the target node to the source node. In complex contagion processes, cycles act as an amplifier. In social networks, this amplification effect could be interpreted as an echo chamber for ideas.
We have also seen how in-centrality and out-centrality characterize the nodes of the network when everything else is assumed to be equal, i.e. when the edges of the network are undirected and when the edge probabilities are the same for all edges of the network. In this case, the network topology determines the in-centrality and out-centrality values of the nodes. We have seen how peripheral nodes tend to be the least influential nodes (lower out-centrality values), whereas more central nodes have more influence (higher out-centrality values). This phenomenon seems to persist as edge weight increases. We have also seen how, for low edge probabilities, peripheral nodes are influenced the least due to lower in-centrality values, whereas more central nodes are being influenced the most because of higher in-centrality values. However, as the edge weight increases, at some point the situation reverses itself, and peripheral nodes become the most influenced nodes, and the most central nodes become the least influenced nodes. In real social networks, for example, it is plausible to think that the peripheral nodes are the least influential nodes, whereas the most central nodes are the most influential nodes. Additionally, in real social networks, nodes can be so isolated that they are influenced very little by the rest of the network due to their isolation, whereas more central nodes are influenced by at least their immediate neighborhood. The in-centrality and out-centrality defined by the influence spreading model produce behavior that aligns well with the aforementioned behavior of a real social network.
The proposed ISM betweenness centrality has been shown to be capable of revealing what standard betweenness centrality cannot. For example, the ISM betweenness can highlight the importance of alternative routes in a network. It also emphasizes nodes that act as important bridges in a network. However, the ISM betweenness centrality has a great computational cost. Calculating the ISM betweenness centrality requires that the influence spreading algorithm be run again for every single node, since removing a different node from the network creates a new network. The ISM betweenness centrality reveals more of the network topology than the standard betweenness centrality, but this advantage comes at a greater computational cost.
8 Conclusion
Many standard centrality measures, such as closeness and betweenness centrality, are based on shortest paths in the network. The influence spreading model defines centrality metrics based on unrestricted paths, walks, and cycles. We have seen how there are similarities between the centrality metrics defined by the influence spreading model and standard centrality metrics, especially for low edge probabilities. We have also seen how the centrality metrics defined by the influence spreading model can reveal information that centrality metrics based on shortest paths cannot. The novelty presented in this article is the detailed examination of the behavior of ISM betweenness centrality, out-centrality, and in-centrality in comparison to standard centrality metrics.
We have demonstrated that ISM betweenness centrality can reveal the importance of alternative routes while maintaining similarity to standard betweenness centrality. We have also demonstrated that out-centrality can detect central nodes as influential nodes and in-centrality can detect peripheral nodes as the most influenced nodes when all else is equal (network edges are undirected and all edge probabilities are equal). These results give validity to out-centrality and in-centrality measures, since we achieve results that are logical. Therefore, we can infer that when studied networks are made more realistic with directed edges and different edge probabilities, we will see even clearer differences between the roles of nodes when our influence-based centrality metrics are used.
Appendix A Appendix
The influence spreading model uses probabilities as edge weights. However, standard betweenness centrality and closeness centrality require a distance to be defined between nodes. In real-world networks, there is some sort of relationship between the influence spreading probability and the distance between nodes. For example, during the COVID-19 pandemic, social distancing was encouraged in order to reduce the probability of infection. Therefore, there is a need to map edge probabilities to edge distances in a systematic way. This is especially true if the edge probabilities are not the same for all edges.
One possible interpretation of a shortest path in a graph which has probabilities as edge weights is to interpret the path with the highest probability as if it were the shortest path. If we denote $p_{1},p_{2},\dots,p_{m}$ as the edge probabilities in a graph $G=(V,E)$, where $V$ is the set of nodes, $E$ is the set of edges between nodes, and $m=|E|$ is the number of edges, then our task is to find a path between source node $s$ and target node $t$ that maximizes the product $p_{(s,a)}\cdots p_{(b,t)}$, where $(s,a),\dots,(b,t)\in E$ are the edges and $s,a,\ldots,b,t\in V$ are the nodes on the path that maximizes the probability of influence between nodes $s$ and $t$.
Maximizing the product $p_{(s,a)}\cdots p_{(b,t)}$ is equivalent to maximizing the logarithmic sum logp_(s, a) + …+ logp_(b, t) which is equivalent to minimizing -logp_(s, a) - …- logp_(b, t) which is again equivalent to minimizing log1p(s, a) + …+ log1p(b, t). In the previous formulation, we assumed that the shortest path contains at least one intermediary node. If that is not the case, then naturally the shortest path between $s$ and $t$ is the edge $(s,t)$.
When $0<p_{i}\leq 1$ for all $1\leq i\leq m$, then $\log(1/p_{i})\geq 0$ for all $1\leq i\leq m$. Now, $\log(1/p_{i})$, for all $1\leq i\leq m$, can be used as a nonnegative distance between nodes, because as probability $p$ increases the distance $\log(1/p)$ decreases, i.e. $\log(1/p)$ is a strictly decreasing function of $p$. Furthermore, we can now use Dijkstra’s algorithm for finding shortest paths in the network, since edge weights are non-negative [thomas2009]. In this case, Dijkstra’s algorithm picks the edges with the lowest distance $\log(1/p_{i})$. Therefore, with this transformation the path $i$ with the lowest value $\log(1/p_{i})$ is chosen i.e. the highest probability path is chosen for the shortest path. In the special case, where $p_{i}=1$ for all $1\leq i\leq m$, then the problem reduces to finding the path with the fewest edges.